A 20-kg hoop travels around a 0.50-m radius circle with an angular velocity of 10 rad/s. The magnitude of its angular momentum about the center of the circle is:


1
Expert's answer
2020-11-25T10:50:25-0500

The angular momentum of a particle about the center of rotation is given by the following formula:


"L = \omega mr^2"

where "m = 20kg" is the mass, "r = 0.5m" is the radius of the circle and "\omega = 10\space rad/s" is the angular velocity. Thus, obtain:


"L = \omega mr^2 = 10\cdot 20\cdot 0.5^2 = 50\space m^2\cdot kg/s"

Answer. "50\space m^2\cdot kg/s".